OpenStudy (anonymous):

solve the initial value problem: (1/x)dy/dx=(ysinx)/(y^2+1), y(pi)=1 I separate both sides, then I integrate which leaves me with, 1/2y^2+lny=-cosx+c but I am lost after this. I've tried reading the book, but it was no help. Please I really need help.

myininaya (myininaya):

hmm how did you get the other side (the right hand side)

myininaya (myininaya):

you did integrate x *sin(x)?

myininaya (myininaya):

by separation of variables you did get: \[\frac{y^2+1}{y} dy =x sin(x) dx \]right ?

OpenStudy (anonymous):

no I did not get xsin(x)dx, I got sin(x)dx

myininaya (myininaya):

well don't you have 1/x on that one side?

myininaya (myininaya):

\[\frac{1}{x} \frac{dy}{dx}=\frac{y \sin(x)}{y^2+1} \] multiply x dx on both sides giving you \[dy=\frac{y \sin(x)}{y^2+1} x dx\] multiply both sides by (y^2+1)/y

OpenStudy (anonymous):

Ah , I suppose that is where I messed up at.

myininaya (myininaya):

Let me know when you have fixed that part and I will help you with the rest or this part if you need further help on this particular part

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

So now I get 1/2y^2+lny=-xcosx+sinx+c I know that I am supposed to solve for y so that I can set the equation equal to the initial condition of y(pi)=1, but I do not know how to solve for y in this problem.

myininaya (myininaya):

a lot better I just want to make one change \[\frac{y^2}{2}+\ln|y|=-xcos(x)+\sin(x)+C\] you don't have to solve for y now replace x with pi and y with 1 because y(pi)=1 says so

OpenStudy (anonymous):

oh, ok so with that then I get c=1/2+1

myininaya (myininaya):

hmmm...doesn't seem right

myininaya (myininaya):

so you replaced x with pi and y with 1 so you have \[\frac{1}{2}+0=-\pi \cos(\pi)+C\]

myininaya (myininaya):

and of course cos(pi) is -1

myininaya (myininaya):

but don't forget that is being multiplied by -1

OpenStudy (anonymous):

ok I see I forgot about the pi in front of -cosx again. now I get c=1/2-pi

myininaya (myininaya):

agreed that is good stuff replace that c above with 1/2-pi and you are done

OpenStudy (anonymous):

oh man so many mistakes I made. Thank you.

myininaya (myininaya):

i only solve for y when it is feasible

myininaya (myininaya):

i think you did okay

OpenStudy (anonymous):

plus I didn't know I could plug in the 1 for y that was a big help

myininaya (myininaya):

yep y(A)=B mean replace x (or t) with A and replace y with B

myininaya (myininaya):

I said or t because t is commonly used too